OpenStudy (anonymous):

I need to know how to solve 3i^2+14i+24 / 9i^2+24i+16

OpenStudy (dls):

\[\LARGE \frac{ 3i^2+14i+24 }{ 9i^2+24i+16}= \frac{21+14i}{7+24i}=1+ \frac{14}{7+24i}\] You can solve further by rationalizing the other part.

OpenStudy (dls):

I used the fact that \(\LARGE i^2=-1\)